OpenStudy (anonymous):

YZ has endpoints y(-1,11) and Z(-1,-3) find the length

OpenStudy (anonymous):

Use the distance formula. Use google to look it up.

OpenStudy (across):

You should know the distance formula:\[D=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]and use it:\[\sqrt{(-1+1)^2+(-3-11)^2}=\sqrt{(-14)^2}=14\]But notice how in this case both points have the same x-intercept. In that case, all you have to do is add/subtract accordingly.

OpenStudy (anonymous):

The midpoint of GH is m(5,3) one is G(1,-5) Find the other coordinate of H